1a. What are the genotype frequencies and allele frequencies in a population with 20 AA genotypes, 25 AB genotypes, and 5 BB genotypes?

1b. What is the chance that you are a carrier for a given trait, if allele "A" is 20% and allele "a" is 80%?
Answer #1

Answer:

1a).

Observed genotype frequencies:

The AA genotype frequency = 20/50 = 0.4

The AB genotype frequency = 25/50 = 0.5

The BB genotype frequency = 5/50 = 0.1

Allele frequencies:

20 AA = 20+20 = 40 A

25 AB = 25 A & 25B

5 BB = 5+5=10 B

Total alleles = 40 + 50 + 10 = 100

Total "A" alleles = 40+25 = 65

Total "B" alleles = 25+10 = 35

Frequency of allele "A" = 65/100 = 0.65

Frequency of allele "B" = 35/100 = 0.35

Expected genotype frequencies:

The AA genotype frequency = 0.65 * 0.65 = 0.4225

The Ab genotype frequency = 2 * 0.65 * 0.35 = 0.455

The BB genotype frequency = 0.35 * 0.35 = 0.1225

1b).

A = 20% = 0.2

a = 80% = 0.8

The carrier frequency = 2pq = 2Aa = 2 * 0.2 * 0.8 = 0.32